Chemistry

The fifth is a unit of measurement equal to a volume of 1/5 gallon. In the late nineteenth century bottles of distilled liquor were bought in units of a fifth. Today, liquor is sold in bottles measured in mL. Given that there are  and , how many milliliters are there in ?
Answer #1

Ans. 1 fifth = 1/5 gallon = 0.2 gallon

                                    = 0.2 gallon x (4 qt / gallon)

                                    = 0.8 qt

                                    = 0.8 qt x (0.946 L / qt)

                                    = 0.7568 L

                                    = 0.7568 L x (1000 mL/ L)

                                    = 756.8 mL

Therefore, 1 fifth = 756.8 mL

# Now,

            1 mL = (1/756.8) fifth = 0.001321 fifth

Therefore, 1 mL = 0.001 fifth